Grade 11Organic Chemistry

Find mass of urea required to make an 50gm aqueous solution in which the mole fraction of urea is 0.077

Nishita
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Molar mass of urea = 60 gm / mol -150 gm  aqueous soln of urea means = 1000 gm of water contains 50 gm of urea,  so total weight = (1000+50)= 1050 gm   Mass = mole fraction  = mass / total weight ,Substituting  the values = 0.77 = m / 1050 = m = 0.77 x 1050 = 808.5 gm